# Privacy by Design for Self-Sovereign Identity Systems: An in-depth Component Analysis completed by a Design Assistance Dashboard

### v1

@@ Abstract
The purpose of this article is to help SSI solution designers make informed choices to ensure that the designed solution is privacy-friendly. The observation is that the range of possible solutions is very broad, from DID and DID resolution methods to verifiable credential types, publicly available information (e.g. in a blockchain), type of infrastructure, etc. As a result, the article proposes (1) to group the elementary building blocks of a SSI system into 5 structuring layers, (2) to analyze for each layer the privacy implications of using the chosen building block, and (3) to provide a design assistance dashboard that gives the complete picture of the SSI, and shows the interdependencies between architectural choices and technical building blocks, allowing designers to make informed choices and graphically achieve a SSI solution that meets their need for privacy.
@@ 1 Introduction
Although the SSI community is converging on standards such as DID (Decentralized Identifier), VC (Verifiable Credential) and others, in terms of design and implementation, SSI solutions couldn’t be more different.
@@
Based on these observations, this paper is intended to help designers and developers of SSI solutions to gain a clearer understanding of the SSI building blocks (hereafter referred to as components) and their functional interdependencies, to understand the privacy implications of certain SSI technological and architectural choices, and to make informed decisions to ensure that the designed solution is privacy-preserving.
@@
A four (4)-layer model for abstracting a SSI system, complemented by the identification of the set of privacy properties most relevant to each layer.
@@
An in-depth privacy analysis of each SSI component is provided to understand the privacy implications of using the selected component.
3.
Identification of the SSI building blocks interdependency. This interdependency limits which components from a given layer are compatible with which components from other layers.
4.
A novel classification of wallet applications based on their architectures.
5.
@@
A design assistance dashboard, which is a graphical representation of the model, called a design graph. It consists of displaying a set of interconnected components belonging to different layers. It helps to show dependencies between SSI layers and components, and conflicts between components. The design graph can be integrated into the computer-aided design of SSI systems. By providing a quick view of SSI systems, it enables the SSI system designer to quickly analyze SSI system proposals and make an informed choice for a SSI solution that satisfies the need for privacy.
@@
The remainder of this paper is organized as follows. Section 2 provides the background on SSI systems, introduces our 4-layer model of SSI systems with the SSI components in each layer, and defines the privacy requirements along with a clear positioning of the privacy requirements per SSI layer. Section 3 presents the related work on the SSI layered models and some privacy related work, highlighting the lack of granularity in existing work to analyze the privacy of SSI systems. The next four sections review the four layers of the model, highlighting the diversity of possible approaches and the privacy implications of each component. 4 analyzes and compares the possible infrastructures to be used with SSI, the base layer of SSI systems. Section 5 addresses the identifiers and cryptographic material needed for identification and authentication in SSI systems, focusing on the most prominent and statistically used SSI identifiers. Section 6 examines SSI credentials, the critical presentation phase of SSI credentials by their owners, and revocation, which is a mandatory feature but often conflicts with the need for privacy. In Section 7, the upper layer of the SSI systems that make up the wallet application is presented with a novel classification of wallets based on their architectures. After presenting our original design assistance dashboard in Section 8, a conclusion is provided in Section 9. The paper also provides five appendices that allow the reader to go further in the privacy analysis of SSI systems.
2 Background on SSI Systems and Privacy
@@ 2.1 SSI Systems
These roles are interchangeable, meaning that, depending on the context, a user may well take on one or other of these roles. Interaction between the different stakeholders is made possible by several SSI building blocks that support the generation, presentation, and verification of credentials. The layered representation of these defined building blocks is illustrated in Figure 2 (cf. Section 3 for arguments in favor of this model) and refers to the following layers:
@@ 2.2 Privacy Requirements in SSI
Referring to the SSI layered model, it is important to understand that the slightest lack of privacy in one layer calls into question the privacy of the whole. In the same vein, a privacy requirement satisfied in one given layer becomes an inherited feature for other layers. For example, if the Identifiers and Cryptographic material layer satisfies non-correlation or pseudonymity by providing identifiers like link secrets111Link secrets are similar to master secret keys used in IDEMIX [16]. They enable Anonymous Credential systems to bet set up, which are the basis for widely used credentials such as AnonCreds., the Credentials and Presentations layer will be able to build upon it, using privacy-preserving signature schemes, to provide anonymous credential systems, which will then satisfy additional privacy requirements, such as selective disclosure with ZKP and credential issue/multi show unlinkability. This requires building up the design in a coherent way in which every layer makes maximum use of privacy mechanisms implemented in other layers. Figure 3 is a reworking of Figure 2 in which we map the privacy properties to the different SSI layers.
SSI systems must satisfy these properties according to the level of authentication needed by each verifier. For instance, proving to an online service provider that a holder is over 18 can be done by showing a government issued ID (like taking a passport photo) that contains the date of birth. However, this means that under several data protection laws (like GDPR), the online service provider is now a collector and/or a processor of personal data. Instead, the selective disclosure property with ZKP and specific anonymous credential, e.g. AnonCreds, can be implemented so that the holder can prove he or she is over 18 thanks to a government-issued credential, while not revealing any other information.
@@ 3 Related Work
4 Infrastructure
The SSI infrastructure is used for one or more of the following reasons:
@@ 4 Infrastructure
Both blockchain and web-based technology can operate as an open or closed infrastructure, which affects the properties that can be expected of them. Open infrastructure, which are operated by any entity on a voluntary basis, are transparent and accessible to everyone. They are harder to protect from curious actors which may attempt to undermine the privacy of users’ activities (non-correlation property) and from malicious actors, which can try to corrupt the working of the infrastructure itself. A contrario, closed infrastructures that are owned and controlled by a few entities, usually have the advantage of providing a higher level of trust compared to closed infrastructures. However, they raise more concerns about privacy, possible illegal processing of user data, lack of transparency and lack of user autonomy.
@@ 4.1 Blockchain infrastructures
Public blockchains do not promote access control or other measures to limit access to data. On the contrary, they promote full transparency and openness, meaning that anyone can be part of the network and see the data on the ledger. Transparency and openness are not necessarily bad for privacy, but they make certain use cases where it is necessary to control who can access the data, such as private transactions between two entities, infeasible on public blockchains. It is still possible to conduct private transactions off-blockchain, but this means missing out on the formalized way of conducting transactions on a blockchain and the decentralized trust provided by the platform. This would require a level of data segregation (separatiing data into silos and controlling access to it on the infrastructure) that not all blockchains or DLTs can provide. Fortunately, newer generations of blockchain platforms integrate access control and permissions to manage who can be part of the blockchain, who can access what data, and even provide private sub-ledgers maintained by subsets of blockchain members. Hyperledger Fabric, with its membership service providers (MSP) and private channels, is a good example.
@@ 4.3 Comparing Web-based and Blockchain Infrastructures for SSI
5 Identifiers and Cryptographic Material
@@ 7 Wallet Applications
We distinguish two main types of digital wallets: hot wallets, which are online wallet services hosted in the cloud, and cold wallets, which are installed on the user’s device with local storage for keys and credentials. SSI wallets are generally cold wallets, as the applications come with a local component called a edge agent, which is installed on the user’s device, and an online component, called a cloud agent or mediator, that acts as a gateway between the wallet and the SSI infrastructure and as a mediator between different edge agents. Separating the cloud agent from the edge agent is an important logical procedure, as it allows us to better understand the privacy of a wallet application (cf. Figures 7 and 7).
PS: Some, like the cryptocurrency wallets communities, consider any internet-connected device to be a hot wallet. In this paper, we think of digital wallets as a spectre that goes from hot - fully online - to cold - fully offline.
@@ 7.1 Hot Wallets and Cold Wallets
In the middle of the spectre shown in Figure 8, we find wallets with an online component and an offline component, e.g.. Metamask wallet [59] which runs as a browser-extension (more about it below). The left end of the spectre includes wallets that are fully online, hosted by a cloud service, and accessible to users via a username and login. In such cases, all creation and storage and communication functions are performed online. Between Metamask and fully hot wallets with no local storage, we find several examples like Exodus wallet [60], Coinbase wallet [61] and others. To correctly position a wallet solution on the spectre, we need to identify which part of the wallet application (edge/client side - cloud/server side) does what functionality. Also note that the wallet examples we cite in this spectre are not just SSI or identity management wallets, but also include cryptocurrency wallets designed to hold private keys and tokens.
@@ 7.2 Cloud Agents
The cloud agent is the component of the wallet that allows the user to interact remotely with other entities, such as issuers and verifiers and other holders. Apart from cases where the entities can physically interact by scanning QR codes and directly exchanging credentials via NFC or Bluetooth or any kind of local connectivity between devices, a cloud agent is required to exchange credentials and eventually messages between users. The cloud agent is an online server in most cases - except in cases where we rely on a peer-to-peer network to forward messages and credentials between peers, for example. Quiet [62] - which is always available to receive and forward exchanges and transactions between edge agents and the blockchain. It has several functionalities such as implementing resolver APIs to handle DID resolution requests coming from edge agents, providing APIs to read and write to the blockchain such as publishing new DIDs or updating published keys and DID documents, forwarding messages and credentials between edge agents, storing messages when the edge agent is unavailable, etc.
@@
The dependency on the cloud agent is the factor we use to place a wallet solution on the hot/cold spectre in Figure 8. There are many privacy considerations that must be taken into account to determine how much a cloud agent should be involved and depend on, but also performance considerations that sometimes make the need for a cloud agent unavoidable. For example, wallets that are designed with messaging capabilities and remote proofing will need a cloud component to handle these online transactions. Other wallets, such as those designed for in-person credential exchange via QR codes, do not require a cloud agent in most cases. However, what is generally required for wallets is the ability to resolve DIDs to obtain public keys, proofing methods and communication endpoints, and this requires a reliable resolver that must be able to resolve - in several cases - different DID methods. This means Internet connectivity and read access to various blockchains and infrastructures.
@@ 7.4.2 DID-Auth Protocol
8 Design Assistance Dashboard
Figure 10 shows our dashboard display, which takes our layered model and represents several possible design choices in the form of design paths. The purpose of the dashboard is to assist designers in the selection of SSI architectural and technical components to build a complete SSI solution from a privacy perspective. Two design paths in blue and green are presented in the figure.
We also represent the main dependencies between components across layers (not all dependencies are shown for simplicity) and optional design possibilities. Most SSI identifiers have a functional dependency on either blockchain, DLT, or web-based infrastructures. Some DID methods and identifiers do not require infrastructure, such as did:key and link secrets. Credentials also depend on the identifiers used to bind credentials to their issuers, their holders and the associated keys used to sign such credentials. They also depend on the infrastructure used to publish the credential schema and definitions, or even to publish credential proofs or hashes. Finally they depend on the wallet and its protocols for storage and exchange. The wallet relies on the communication protocols, both between wallets (wallet to wallet) and between wallet components themselves (edge to cloud).
@@ 8 Design Assistance Dashboard
Figure 10 is not an exhaustive list of dependencies and design paths, but rather a visual representation of SSI design dependencies and paths. The diagram is used by selecting different components from each layer and linking them together to form a SSI system. Dependencies are not optional links, i.e. if a component is selected from one layer and has a dependency on a component from another layer, then both must be selected. Note that multiple components of the same layer can be selected in an SSI system. For example, a system may have multiple types of identifiers, use different types of infrastructure, and offer different types of credentials.
@@
To introduce a privacy indicator in Figure 10, we refer to our earlier analysis of different components in relation to each other. In line with the privacy requirements a component should provide in a layer, we identify three (3) levels of privacy ranging from low, medium, to good. This privacy level only indicates the component’s ability to provide the privacy features that have been mapped to the layer in which the component operates. Note that certain components - such as did:web - may present a low level of privacy as an identifier when used for a normal user-holder, but are still relevant because they are intended to be used for public actors, such as issuers and verifiers, and do not need to meet the privacy requirements mapped to their layer.
@@
Path 1 and Path 2 (in blue and green) in Figure 10 are suggested for analysis131313These are examples of possible design paths illustrated on Figure 10, they are not recommendations and do not represent the best privacy-preserving design choices.:
Path 1 (blue): We choose to use the did:web method, which means that we need a web-based infrastructure and DNS. We use did:web identifiers to identify and authenticate issuers that provide and sign credentials using the keys and endpoint specified in the DID document, which is easily retrieved from the issuer’s server. The security and privacy of this solution depends on the DNS protocol and servers. Choosing did:web does not limit the keys we can associate with the identifier and specify in the DID document, which means maximum flexibility in choosing the VC scheme and signature scheme later. The VC schema and definition should also be published as resources on common web-based registries or even stored as web resources like the DID documents. JWT are selected as the VC and stored on the cloud agent. Storage in the cloud agent means no recovery issues, easy access from any device, and faster exchange through efficient routing. Holders can be identified using did:key identifiers that have direct resolution and no dependencies. DID-Auth is used to authenticate actors via QR codes, and for simplicity, JWT may also be displayed as QR codes to be scanned by verifiers. Note that the resolution of did:web can optionally use a VPN or pass through the universal resolver, which improves the privacy aspects of the identifier resolution and limits activity tracing or correlation of the resolution by the owner of the server on which the DID document is stored (cf. Section 5.2).
Path 2 (green): In this design path, we choose a very common identifier method, the did:indy method, which requires a (decentralized) Verifiable Data Registry to publish identifiers and DID documents. The did:indy identifiers will be used by issuers, verifiers and as main public identity for holders. AnonCreds [41] are the VC of choice in this example, so we also use another identifier, link secrets, to bind credentials to their holders in a privacy preserving manner. Hyperledger Indy is the blockchain of choice that did:indy identifiers generally rely on. Hyperledger Indy also provides storage for revocation lists and VC schemas and definitions. We chose to store credentials on the edge agent and exchange them as messages over the DID-Comm protocol for later verification by service providers or relying parties. DID-Comm allows messages to be exchanged in a simplex way and can turn the edge agent into a peer that is part of a peer-to-peer exchange network that does not require web servers. Intermediate peers require little trust because DID-Comm encryption provides end-to-end privacy and security, and credentials can be exchanged between peers without intervention.
9 Conclusions
Given the importance of performant and privacy-preserving SSI systems, especially in the current context where most countries are deploying national digital identity solutions, it is important for software vendors to prioritize offering privacy by design in their systems to maximize compliance with privacy legislation. The sheer number of components that make up a SSI solution, and the number of possible combinations, make it difficult to find one’s way around and design a correct SSI system, even more so if privacy is taken into account. This article takes an in-depth look at each component, covering the cryptographic material and identifiers, the credential design and signature schemes, the wallet application design, the inter-wallet communication protocols, and the underlying infrastructure.
This paper covers the design choices in SSI systems, along with a careful decomposition of SSI systems into layers to highlight a more generic architecture that combines different components and technologies used for each functionality.
Through this dissection and understanding of SSI systems, we have been able to show how different architectural and technological choices affect the security, privacy, and performance of SSI solutions. We also provide a design assistance dashboard that graphically represents several SSI components in their respective layers, and shows the dependency and functional relationship between such layers and components. The dashboard allows a designer to understand the privacy level of each SSI component thanks to a color code, to assemble SSI components to achieve a SSI solution, and to quickly assess the resulting privacy level of that solution.

### v2

# A Systematic Review and Layered Framework for Privacy-by-Design in Self-Sovereign Identity Systems

@@ Abstract
The design of SSI systems is complex, often characterized by a large number of components and architectural choices because the current SSI communities differ on how to create identifiers, how to build and present credentials, and even how to design a user wallet. SSI stacks developed by different organizations provide different privacy features for different privacy needs.
This paper performs a systematic mapping and review of SSI components and technologies into a novel four-layer privacy framework to address the design complexity of SSI systems. Based on this review, we provide an accompanying Design Assistance Dashboard (DAD). The DAD shows the interdependencies between SSI components in different layers, and maps these components to different privacy requirements and considerations, even providing a simple privacy class for each component.
@@ 1 Introduction
Although the SSI community is converging on standards such as DID (Decentralized Identifier), VC (Verifiable Credential) and others, in terms of design and implementation, SSI solutions could not be more different.
@@
Based on these observations, this paper is intended to help gain a deep understanding of the SSI building blocks (hereafter referred to as components) and their functional interdependencies, and to understand the privacy implications of possible architectural choices.
@@
A deep survey and review of SSI ecosystems. The major SSI components are mapped into a layered framework that associates technological components with privacy requirements and considerations.
@@
A Design Assistance Dashboard (DAD), which is a graphical representation of the framework, known as a design graph. It consists of displaying a set of interconnected components belonging to different layers. It helps to show dependencies between SSI layers and components, and conflicts between components. The design graph can be integrated into the computer-aided design of SSI systems. By providing a quick view of SSI systems, it enables the SSI system designer to quickly analyze SSI system proposals and make an informed choice for a SSI solution that satisfies the need for privacy.
@@
The remainder of this paper is organized as follows. In Section 2, we present the research methodology and describe how the SSI ecosystems were mapped and reviewed. We also explain how the privacy-oriented framework and DAD were constructed. Section 3 provides an overview of SSI systems, introduces our four layer privacy model of SSI systems. Section 4 presents the related work on the SSI layered models and some privacy related work, highlighting the lack of granularity in the existing work to analyze the privacy of SSI systems. The next four sections review the four layers and their components, highlighting the diversity of possible approaches and the privacy implications of each component. 5 analyzes and compares the possible infrastructures to be used with SSI, the base layer of SSI systems. Section 6 addresses the identifiers and cryptographic material needed for identification and authentication in SSI systems, focusing on the most prominent and statistically used SSI identifiers. Section 7 examines SSI credentials, the critical presentation phase of SSI credentials by their owners, and revocation, which is a mandatory feature but often conflicts with the need for privacy. In Section 8, the upper layer of the SSI systems that make up the wallet application is presented with a novel classification of wallets based on their architectures. After presenting our original Design Assistance Dashboard (DAD) accompanied by a step-by-step application guide in Section 9, Section 10 is dedicated to discussion, challenges and future works.
2 Methodology
In order to construct a framework that is comprehensive in its review and analysis, yet practical in its application, we adopted a two-phase methodology. In the first phase, we conducted a thorough technology and literature review to map and analyze SSI components and the ecosystem. In the second phase, we synthesized our findings to create our proposed layered privacy framework and the accompanying Design Assistance Dashboard (DAD).
Phase 1: Systematic Review
This phase is crucial for deconstructing the complex SSI landscape, which ranges from new standards (both accepted and still in draft form) to SSI working groups under different organisations, literature and academic proposals closely related to SSI, and industrial products such as SSI wallets provided by SSI vendors. We have a multi-source review process:
1. Analysis of standards and specifications: An in-depth analysis of current SSI standards and specifications, ranging from adopted standards to simple editor drafts. Technical documentation from the W3C on core data models such as DIDs, VCs and VPs was studied and analysed extensively. Protocols from DIF, such as DID-Comm (v1 and v2), and early DID-Auth proposals from RWOT (Rebooting the Web of Trust), were also studied. This enabled us to develop a foundational understanding of how SSI components are designed to function.
2. Review of open-source implementations: The gap between specifications and practice is bridged by reviewing the architectures and documentation of influential open-source SSI stacks and projects. The most well-known example is the Linux Foundation’s Hyperledger ecosystem, which comprises Hyperledger Indy, Aries and AnonCreds. This in-depth analysis sheds light on how theoretical standards are implemented in real-world software stacks and the trade-offs between security, privacy, and performance.
3. Assessment of real-world products and industry data: To ensure that our framework reflects currently used technologies, we prioritised the SSI components with the widest adoption. This selection was guided by industry data, such as the CheckD survey of 37 SSI vendors (cited in checkD-survey), which provides quantitative data on the market share of competing DID methods and VC schema. We also closely studied some SSI wallets and even exchanged views with the technical teams behind them to understand the practical challenges and implementation patterns.
Phase 2: Framework Synthesis and Development
In this phase we synthesized the findings from phase 1 in two steps:
1. Construction of the layered privacy framework: We categorized the identified components based on their functional role in a typical SSI interaction (e.g., identification, attestation and user interaction). This bottom-up, functional grouping naturally led to the development of our proposed four-layer model: Infrastructure, Identifiers & Cryptographic Material, Credentials & Presentations, and Wallet Applications. While there are similar models (cf. Section 4), we take it a step further by identifying privacy requirements from different sources, such as regulation, industry standards, community discussions. We then map these privacy properties to the four SSI layers. The result is a layered privacy model in which we can analyze each layer and its components from a privacy perspective with defined goals and requirements.
2. DAD creation: DAD builds on the previous layered model, adding two extra steps. Firstly, DAD identifies the dependencies and architectural possibilities between the different components. The second step involves attributing a privacy indicator to each component, which allows the privacy level to be visualized using a simple colour code. Please note that this privacy indicator is a simplified privacy rating based on our analysis and review provided for each component in Phase 1.
DAD is a graphical representation of our SSI review process. It allows users to analyze and design SSI systems by following a simple flowchart on a graphical interface.
This two-phase methodology ensures that our review is not just theoretical, but has a practical application in the form of the DAD.
3 SSI Systems and Privacy
@@ 3.1 SSI Systems
These roles are interchangeable, meaning that, depending on the context, a user may well take on one or other of these roles. The interaction between different stakeholders is made possible by several SSI building blocks that support the generation, presentation, and verification of credentials. The layered representation of these defined building blocks is illustrated in Figure 2 (cf. Section 4 for arguments in favor of this model) and refers to the following layers:
@@ 3.2 Privacy Requirements in SSI
Referring to the SSI layered model, it is important to understand that the slightest lack of privacy in one layer calls into question the privacy of the whole. In the same vein, a privacy requirement satisfied in one given layer becomes an inherited feature for other layers. For example, if the Identifiers and Cryptographic material layer satisfies non-correlation or pseudonymity by providing identifiers like link secrets111Link secrets are similar to master secret keys used in IDEMIX [16]. They enable Anonymous Credential systems to bet set up, which are the basis for widely used credentials such as AnonCreds., the Credentials and Presentations layer will be able to build upon it, using privacy-preserving signature schemes, to provide anonymous credential systems, which will then satisfy additional privacy requirements, such as selective disclosure with ZKP and credential issue/multi show unlinkability. This requires building up the design in a coherent way in which every layer makes maximum use of privacy mechanisms implemented in other layers. Figure 2 also maps privacy requirements to SSI layers.
@@ 4 Related Work
5 SS Infrastructure
An SSI infrastructure is used for one or more of the following reasons:
@@ 5 SS Infrastructure
Both blockchain and web-based technology can operate as an open or closed infrastructure, which affects the properties that can be expected of them. Open infrastructure, which are operated by any entity on a voluntary basis, are transparent and accessible to everyone. They are harder to protect from curious actors which may attempt to undermine the privacy of users’ activities (non-correlation property) and from malicious actors, which can try to corrupt the working of the infrastructure itself. A contrario, closed infrastructures that are owned and controlled by a few entities, usually have the advantage of providing a higher level of trust compared to open infrastructures. However, they raise more concerns about privacy, possible illegal processing of user data, lack of transparency and lack of user autonomy.
@@ 5.1 Blockchain infrastructures
Public blockchains do not promote access control or other measures to limit access to data. On the contrary, they promote full transparency and openness, meaning that anyone can be part of the network and see the data on the ledger. Transparency and openness are not necessarily bad for privacy, but they make certain use cases where it is necessary to control who can access the data, such as private transactions between two entities, infeasible on public blockchains. It is still possible to conduct private transactions off-blockchain, but this means missing out on the formalized way of conducting transactions on a blockchain and the decentralized trust provided by the platform. This would require a level of data segregation (separating data into silos and controlling access to it on the infrastructure) that not all blockchains or DLTs can provide. Fortunately, newer generations of blockchain platforms integrate access control and permissions to manage who can be part of the blockchain, who can access what data, and even provide private sub-ledgers maintained by subsets of blockchain members. Hyperledger Fabric, with its membership service providers (MSP) and private channels, is a good example.
@@ 5.3 Comparing Web-based and Blockchain Infrastructures for SSI
6 SSI Identifiers and Cryptographic Material
@@ 8 SSI Wallet Applications
We distinguish two main types of digital wallets: hot wallets, which are online wallet services hosted in the cloud, and cold wallets, which are installed on the user’s device with local storage for keys and credentials. SSI wallets are generally cold wallets, as the applications come with a local component called a edge agent, which is installed on the user’s device, and an online component, called a cloud agent or mediator, that acts as a gateway between the wallet and the SSI infrastructure and as a mediator between different edge agents. Separating the cloud agent from the edge agent is an important logical procedure, as it allows us to better understand the privacy of a wallet application (cf. Figures 6 and 6).
PS: Some, like the cryptocurrency wallets communities, consider any internet-connected device to be a hot wallet. In this paper, we think of digital wallets as a spectrum that goes from hot - fully online - to cold - fully offline.
@@ 8.1 Hot Wallets and Cold Wallets
In the middle of the spectrum shown in Figure 7, we find wallets with an online component and an offline component, e.g.. Metamask wallet [59] which runs as a browser-extension (more about it below). The left end includes wallets that are fully online, hosted by a cloud service, and accessible to users via a username and login. In such cases, all creation and storage and communication functions are performed online. Between Metamask and fully hot wallets with no local storage, we find several examples like Exodus wallet [60], Coinbase wallet [61] and others. To correctly position a wallet, we need to identify which part of the wallet application (edge/client side - cloud/server side) does what functionality.
@@ 8.2 Cloud Agents
The cloud agent is the component of the wallet that allows the user to interact remotely with other entities, such as issuers and verifiers and other holders. Apart from cases where the entities can physically interact by scanning QR codes and directly exchanging credentials via NFC or Bluetooth or any kind of local connectivity between devices, a cloud agent is required to exchange credentials and eventually messages between users. A cloud agent has several functionalities such as implementing resolver APIs to handle DID resolution requests coming from edge agents, providing APIs to read and write to the blockchain such as publishing new DIDs or updating published keys and DID documents, forwarding messages and credentials between edge agents, storing messages when the edge agent is unavailable, etc.
@@
The dependency on the cloud agent is the factor we use to place a wallet solution on the hot/cold spectrum in Figure 7. There are many privacy considerations that must be taken into account to determine how much a cloud agent should be involved and depend on, but also performance considerations that sometimes make the need for a cloud agent unavoidable. For example, wallets that are designed with messaging capabilities and remote proofing will need a cloud component to handle these online transactions. Other wallets, such as those designed for in-person credential exchange via QR codes, do not require a cloud agent in most cases. However, what is generally required for wallets is the ability to resolve DIDs to obtain public keys, proofing methods and communication endpoints, and this requires a reliable resolver that must be able to resolve - in several cases - different DID methods. This means Internet connectivity and read access to various blockchains and infrastructures.
@@ 8.4.2 DID-Auth Protocol
Now that we have provided a deep review of SSI ecosystem and components, we move to introduce DAD, our Design Assistance Dashboard that provides a fast reliable way to analyze and evaluate SSI systems in terms of privacy.
9 Design Assistance Dashboard (DAD) Framework
DAD is a practical application of our layered model, providing a review tool that classifies the privacy of different SSI components and their interdependencies. In this section, we present our DAD and how to apply it.
9.1 DAD Definition
The privacy-oriented layered framework can be represented as a design dashboard if different components in each layer are enumerated and associated with other components from other layers. This association either represents a dependency or a possible design choice, so by systematically associated SSI components across layers to each other, we end up with a dashboard that graphically illustrates SSI design process. Figure 9 shows our dashboard display, with two example design paths, marked in blue and green. Rather than providing an exhaustive representation of dependencies and design paths, the intention is to offer a visual representation of SSI design dependencies and paths, as well as a colour scheme that allows the level of confidentiality associated with each component to be identified at a glance.
@@ 9.1 DAD Definition
The dashboard is used by selecting different components from each layer and linking them together to build an SSI system. Dependencies are not optional links, i.e. if a component is selected from one layer and has a dependency on a component from another layer, then both must be selected. Note that multiple components of the same layer can be selected in an SSI system. For example, a system may have multiple types of identifiers, use different types of infrastructure, and offer different types of credentials.
@@
To introduce a privacy indicator in DAD, we refer to our earlier analysis of different components in relation to each other. In line with the privacy requirements a component should provide in a layer, we identify three (3) levels of privacy ranging from low, medium, to good. This privacy level only indicates the component’s ability to provide the privacy features that have been mapped to the layer in which the component operates. Note that certain components - such as did:web - may present a low level of privacy as an identifier when used for a normal user-holder, but are still relevant because they are intended to be used for public actors, such as issuers and verifiers, and do not need to meet the privacy requirements mapped to their layer.
9.2 Applying DAD: A step-by-step guide
@@
Path 1 and Path 2 (in blue and green) illustrated in Figure 9 are the result of applying the step-by-step guide to two use cases presented below. They should not be taken as recommendations and do not represent the best design choices for privacy.
The step-by-step guide is presented in the flowchart in Figure 10. It identifies seven main steps, the very first being the definition of privacy goals and requirements, with the aim of analyzing and designing an SSI system.
The use cases studied for applying the guide are as follows:
Case A: Public University Diploma System (Path 1 in blue)
1. Define Goals: the primary requirements are public verifiability and high integrity. A university is a public issuer that does not require the privacy of its identifier. Students are the holders, and their privacy is secondary to the need for employers to verify diplomas. The target is medium privacy.
2. Select Infrastructure: Given the university’s public nature and existing web presence, a web-based infrastructure is an efficient choice. DAD shows that web-based infrastructures offer good privacy. The university can host the necessary DID document and credential schema as web-resources on their official website.
3. Select Identifiers: did:web is a web-dependent identifier that uses the domain name of the university. It is a low-privacy identifier, but it is ideal for universities as they that don’t need privacy. For holders who want medium privacy and portability, a non-dependent identifier is best, so did:key is a great choice.
4. Select Credential Schema: JWT offers low privacy but are simple and widely compatible with the web environment.
5. Determine Wallet Architecture: to prioritize ease of access for students from any device, the cloud agent is responsible for storing credentials, thereby accepting the trade-off of reduced user-control for greater convenience. The same cloud agent can also perform DID and key generation alongside cryptographic operations.
6. Choose Protocols: DID-Auth and DID-Comm are the two main SSI wallet protocols. However, there are different ways to implement both protocol. For this blue path, we use DID-auth over OIDC to authenticate students with their university in order to request a diploma as a credential. DID-Comm is not needed since the Cloud Agent and the University/Employers can communicate using other web protocols to communicate without relying on DIDs.
7. Evaluate Full Path: this blue path consists of red, orange and green components. While it successfully creates a public, verifiable credential system, it offers limited privacy for the credential holder. This architecture aligns with the initial goal of prioritizing public verifiability over privacy.
Case B: Verifiable Health Credential System (Path 2 in green)
1. Define Goals: the primary requirement is maximum privacy for the holder. Unlinkability and selective disclosure of sensitive information are required, the target is high privacy.
2. Select Infrastructure: avoiding central control or surveillance requires a decentralized infrastructure. Blockchain with medium privacy is selected as a verifiable data registry.
3. Select Identifiers: a Link Secret offers good privacy for holders, while did:indy with medium privacy is ideal for issuers.
4. Select Credential Schema: AnonCreds is a standard VC schema for credentials in the Hyperledger (Indy, Aries) ecosystem. They offer good privacy and allow for ZKP and Selective Disclosure.
5. Determine Wallet Architecture: exclusive holder control over credentials and identifiers, as well as the rest of their health data, is required. An Edge Agent implementing sensitive cryptographic operations (key and identifier generation, signatures) and storing the health data locally is also required.
6. Choose Protocols: using a cloud agent to mediate between edge agents is an option, but peer-to-peer, end-to-end encrypted communication between edge agents is preferable. Both DID-Auth and DID-Comm offer good privacy for authentication and data exchange.
7. Evaluate Full Path: the resulting path consists entirely of green and orange components. It creates a system that prioritizes the privacy, control and unlinkability of the holder, thus meeting the initial goal of providing a high level of privacy.
10 Discussion, Open Challenges and Future Work
In this paper, we conducted an in-depth review of the SSI ecosystem and synthesized our findings into a privacy-oriented SSI layered framework. This framework is accompanied by a Design Assistance Dashboard (DAD), which is a tool that enables the quick analysis of SSI dependencies and assessment of the privacy levels of a given architecture. Unlike other models that focus on the governance or interoperability issues of SSI systems, such as ToIP, our paper presents a layered, privacy-first dissection of SSI that highlights the trade-offs often made in real-world implementations. However, despite the promise and utility of SSI, our review points to the following open challenges:
True user sovereignty: it is difficult to achieve the SSI principles in practice. Even with a privacy-preserving design, user control is still mediated by multiple third-parties, like the wallet vendor (provider), the operating system of the device running the wallet and the device manufacturer. This can undermine user’s authority over their identity, even when privacy guarantees are implemented at the SSI level.
Recovery and usability: managing keys and identifiers poses a significant challenge in terms of usability. The recovery of keys and credentials is one of the largest barriers to mainstream adoption since there is a high risk of permanent loss131313The Bitcoin community is familiar with this, estimates suggest that 11-18% of the total Bitcoin supply ha been permanently lost in wallets with no access due to key loss. https://www.ledger.com/academy/topics/economics-and-regulation/how-many-bitcoin-are-lost-ledger.
Performance cost of privacy: as the credential analysis shows, stronger privacy protection comes at a higher computational cost due to larger key and signature sizes. This performance overhead makes privacy-preserving solutions impractical for large-scale applications or environments with limited resources.
In conclusion, this work provides a foundation for several future works in SSI. Our immediate goal is to continue expanding the DAD by incorporating additional components and highlighting more detailed dependencies and architectural options as new standards and components emerge. The most significant step is to develop the DAD as an interactive digital tool with a graphical user interface (GUI). This would allow users to select SSI components and choose their architecture, with the tool automatically highlighting dependencies, flagging privacy conflicts or redundancies, and potentially calculating an overall privacy score with better-defined metrics, rather than using three classes (good, medium, bad). Such a tool would enhance computer-aided design of SSI and will serve as a design and educational resource for the SSI community. Such a tool would enhance computer-aided design of SSI and serve as a valuable resource for designers and educators within the SSI community.